# Distributional Lusternik–Schnirelmann category of manifolds

### v2

@@ 7. For Alexandrov spaces
We now study A. D. Alexandrov spaces with curvature bounded below. In this section, we will extend some of our results for closed manifolds from the previous sections to a more general class of closed Alexandrov spaces.
7.A. Some definitions
An Alexandrov space is a finite-dimensional, complete, locally compact length space with an inner length metric that locally satisfies a lower curvature bound in the sense of distance comparison. Alexandrov spaces are locally contractible metric ANR [Wu]. We refer to [BGP][BBI][GGNZ, Section 2], and [Mit, Section 6] for their precise definitions. Examples of Alexandrov spaces include orbit spaces of isometric compact Lie group actions on complete Riemannian manifolds whose sectional curvatures are uniformly bounded below. A non-manifold and non-orbifold example is the suspension of ${\mathbb{C}}P^{2}$, denoted $\operatorname{Susp}({\mathbb{C}}P^{2})$. For more examples and constructions, we refer the reader to [GGNZ, Section 2.3].
@@ 7.A. Some definitions
The connected sum of $M$ and $N$ along $(p,q)$, denoted $M\#^{(p,q)}N$, is defined as the space $((M\setminus U)\sqcup(N\setminus V))/\sim$, where $x\sim\psi(x)$ for all $x\in\partial(M\setminus U)$.
Indeed, $M\#^{(p,q)}N$ is a closed Alexandrov $n$-space with curvature bounded below — see, for example, [FGGGH, Section 3] and [GGNZ, Page 9]). Unlike the case of the connected sum of closed manifolds, in general, the sum $M\#^{(p_{1},q_{1})}N$ need not be homeomorphic to $M\#^{(p_{2},q_{2})}N$, even if $M=N$, $p_{1}=q_{1}$, $p_{2}=q_{2}$, and $\varSigma_{p_{1}}=\varSigma_{p_{2}}$, as shown in [FGGGH, Theorem A] for $n=3$. However, when $\varSigma_{p}=\varSigma_{q}=S^{n-1}$ and there is no risk of confusion regarding $p$ and $q$, we will write $M\#N$ to denote the usual connected sum of $M$ and $N$ along closed $n$-discs.
7.B. Extension of some results
@@
Unlike the case of closed manifolds, the existence of fundamental class and the satisfaction of Poincaré duality is very restrictive in the case of closed Alexandrov spaces, so we will need to be careful while extending our results.
Let $M$ and $N$ be two closed Alexandrov $n$-spaces such that $N$ is a manifold. We can choose regular points $p\in M$ and $q\in N$ (both $M$ and $N$ have infinitely many regular points) to form $M\#^{(p,q)}N$, which we will write as $M\#N$ from here onwards. Let ${\mathcal{O}}_{N}$ be the orientation sheaf on $N$. We follow the conventions from Section 4. The fundamental class $[M\#N]_{R}$ always exists for either $R={\mathbb{Z}}$ or $R={\mathbb{Z}}_{2}$ (a ${\mathbb{Z}}$-fundamental class exists when both $M$ and $N$ are ${\mathbb{Z}}$-orientable) and gets mapped to $[N]$ under the homomorphism induced in homology by the map $\phi:M\#N\to N$ that collapses $M$ to a point. So, the following result is obtained.
7.3 Theorem.
Let $M$ and $N$ be closed Alexandrov $n$-spaces. If $N$ is an aspherical manifold, then $\operatorname{\mathsf{dcat}}(M\#N)=\operatorname{\mathsf{cat}}(M\#N)=n$ for sum along regular points.
The proof for $\operatorname{\mathsf{dcat}}(M\#N)\geq n$ is the same as that of Corollary 4.3 and uses Poincaré duality in the manifold $N$ in ${\mathcal{O}}_{N}\otimes{\mathcal{F}}$ coefficients, where ${\mathcal{F}}$ is a $\pi_{1}(N)$-module such that $\kappa_{n}\in H^{n}(N;{\mathcal{F}})$ is the non-zero primary obstruction to a section of ${\mathcal{B}}_{n}(p_{N}):P(N)_{n}\to N$ due to the asphericity of $N$. Since any closed Alexandrov $n$-space $X$ is an ANR, in particular, locally contractible [Wu], we have $\operatorname{\mathsf{cat}}(X)\leq n$ by [CLOT, Theorem 1.7]. Therefore, $\operatorname{\mathsf{dcat}}(M\#N)=\operatorname{\mathsf{cat}}(M\#N)=n$ is obtained.
@@
Any flat $n$-manifold $N$ is an aspherical Alexandrov $n$-manifold. So, $\operatorname{\mathsf{dcat}}(M\#N)=n$ by Theorem 7.3 for any Alexandrov $n$-space $M$. This gives more examples. For $n\geq 4$, let $P^{n-3}$ be an Alexandrov $(n-3)$-space. For any $t\geq 1$, the product #$\#_{i=1}^{t}\operatorname{Susp}({\mathbb{R}}P^{2})\times P^{n-3}$ is a non-manifold Alexandrov $n$-space. So, we get $\operatorname{\mathsf{dcat}}((\#_{i=1}^{t}\operatorname{Susp}({\mathbb{R}}P^{2})\times P^{n-3})\#N)=n$ by Theorem 7.3 for any flat $n$-manifold $N$. In particular, we can take $N=\mathcal{N}^{k}\times T^{m}$ for $n=k+m$, where $T^{m}$ is the $m$-torus and $\mathcal{N}^{k}$ is the generalized Hantzsche–Wendt $k$-manifold discussed in Example 4.8. Furthermore, Theorem 7.3 also recovers the new computation $\operatorname{\mathsf{dcat}}((\mathcal{N}^{k}\times T^{m})\#(\mathcal{N}^{k}\times T^{m}))=k+m$ obtained in Example 4.8.
In the orientable (or ${\mathbb{Z}}$-orientable) case, a proof simpler than that of Theorem 4.1 and Corollary 4.3 can be given for (possibly non-manifold) Alexandrov spaces.
7.5 Theorem.
Let $M$ and $N$ be closed orientable Alexandrov $n$-spaces. If $N$ is aspherical, then $\operatorname{\mathsf{dcat}}(M\#N)=\operatorname{\mathsf{cat}}(M\#N)=n$ for sum along regular points.
Proof.
Since $N$ is a closed orientable aspherical Alexandrov $n$-space, we can use Theorems 2.3 and 7.2 and [Mit, Theorem 1.8] to get $\operatorname{\mathsf{dcat}}(N)=\operatorname{{\rm cd}}(\pi_{1}(N))=n$. So, in view of Proposition 2.4, the primary obstruction $\kappa_{n}\in H^{n}(N;{\mathcal{F}})$ to a section of the fibration ${\mathcal{B}}_{n}(p_{N}):P(N)_{n}\to N$ is non-zero. Due to orientability, $[M\#N]$ and $[N]$ exist such that $\phi_{*}([M\#N])=[N]$ for the collapsing map $\phi:M\#N\to N$. Clearly, $\phi$ induces an isomorphism of the zeroth homology groups in each coefficient system. Therefore, we have
@@
where $\phi^{*}(\kappa_{n})\in H^{n}(M\#N;\phi^{*}({\mathcal{F}}))$ is the primary obstruction to a lift of $\phi$ with respect to ${\mathcal{B}}_{n}(p_{N})$. Since $\dim(M\#N)=n$, it is the only obstruction to such a lift of $\phi$. Here, ${\mathcal{F}}$ is a $\pi_{1}(N)$-module, hence an abelian group. By Poincaré duality in $N$ in ${\mathcal{F}}$ coefficients (see Theorem 7.2), we have $[N]\frown\kappa_{n}\neq 0$. This means $\phi^{*}(\kappa_{n})\neq 0$. So, $\phi$ does not admit a lift with respect to ${\mathcal{B}}_{n}(p_{N})$. Thus, using Propositions 2.5 (2) and 2.6, we get $n\leq\operatorname{\mathsf{dcat}}(\phi)\leq\operatorname{\mathsf{dcat}}(M\#N)\leq\operatorname{\mathsf{cat}}(M\#N)\leq n$. ∎
@@
We note that the orientability assumption is important to prove Theorem 7.5. If $N$ is not orientable in the sense of [HS], then we will be dealing with ${\mathcal{O}}_{N}\otimes{\mathcal{F}}$ coefficients in $N$ for the cap product $[N]\frown\kappa_{n}$. Here, ${\mathcal{O}}_{N}$ is the “orientation bundle” on $N$, see [Mit, Section 5.3]. So, in this case, while we can still proceed along the lines of Theorem 4.1, we cannot apply Theorem 7.2 to get $[N]\frown\kappa_{n}\neq 0$ because in general, ${\mathcal{O}}_{N}\otimes{\mathcal{F}}\neq{\mathcal{F}}$, even if ${\mathcal{O}}_{N}={\mathbb{Z}}_{2}$.
@@
Following the notations of Example 7.4, we note that Theorem 7.5 also recovers the new computation $\operatorname{\mathsf{dcat}}((\mathcal{N}^{k}\times T^{m})\#(\mathcal{N}^{k}\times T^{m}))=k+m$ obtained in Example 4.8 since $\mathcal{N}^{k}$ is orientable for each odd $k\geq 3$.
Theorems 7.3 and 7.5 both partially extend Corollary 4.3 to some non-manifold cases since they include the possibility of $M$ not being a manifold. In Theorem 7.3, we do not require the orientability of $N$ or $M\#N$ but need $N$ to be a manifold. In Theorem 7.5, we need orientability but do not require any space to be a manifold.
Formally, Theorem 7.5 gives new computations of $\operatorname{\mathsf{dcat}}$ for orientable non-manifold cases. Unfortunately, we do not have an example of a closed aspherical Alexandrov $n$-space that is not a manifold. For $n=3$, it has been conjectured in [GGNZ, Conjecture 5.1] that such examples do not exist.
@@ 8. For non-manifold Alexandrov $3$-spaces
We now focus on three-dimensional Alexandrov spaces that have recently been studied extensively in [GGG][GGS][NZ], and [FGGGH], for example. In this section, using some well-known results from Alexandrov geometry (in particular, from the above papers), we shall make some observations regarding the classical and the distributional LS-categories of these spaces.
8.A. Initial observations
@@
Given a closed Alexandrov $3$-space $X$, let $M_{X}$ be the compact non-orientable $3$-manifold with boundary obtained by removing disjoint regular neighborhoods of the finitely many singular points of $X$. So, the boundary of $M_{X}$ is a finite collection of ${\mathbb{R}}P^{2}$. Let $M_{X}=M_{1}\#\cdots\#M_{n}\#l(S^{2}\propto S^{1})$ be the unique prime decomposition [He], where $l\geq 0$ and $S^{2}\propto S^{1}$ denotes the non-orientable $S^{2}$-bundle over $S^{1}$. For each $1\leq i\leq n$, let $\widehat{M_{i}}$ be the Alexandrov $3$-space obtained from $M_{i}$ by capping off its ${\mathbb{R}}P^{2}$ boundary components. Then,
@@
For some $j\leq n$, if $\widehat{M_{j}}$ is orientable and aspherical, or if it is an aspherical manifold, then $\operatorname{\mathsf{dcat}}(X)=3$ in view of Theorems 7.5 and 7.3, respectively. The latter situation is possible when $M_{j}$ is aspherical with no ${\mathbb{R}}P^{2}$ boundary components.
@@
We use the above notations. By our assumption, $l\geq 1$. Thus, we can write $X=Y\#(S^{2}\propto S^{1})$, where $Y=\widehat{M_{1}}\#\cdots\#\widehat{M_{n}}\#(l-1)(S^{2}\propto S^{1})$. Since $S^{2}\times S^{1}$ is the orientable double cover of $S^{2}\propto S^{1}$, the space $Z=Y\#Y\#(S^{2}\times S^{1})$ is a cover of $X$. Since $X$ is a CW complex, so is its cover $Z$. Thus, $(Z,S^{2})$ is a good pair [Ha]. So, we can proceed using the Mayer–Vietoris sequence for cohomology of the pair $(Z,S^{2})$ along the lines of [J2, Remark 8.2] to get $2=c\ell_{{\mathbb{Q}}}(S^{2}\times S^{1})\leq c\ell_{{\mathbb{Q}}}(Z)\leq\operatorname{\mathsf{dcat}}(Z)$. Finally, we use Theorem 2.7 to conclude that $2\leq\operatorname{\mathsf{dcat}}(Z)\leq\operatorname{\mathsf{dcat}}(X)$. ∎
@@
Due to these reasons, the subsequent results in [GLGA] and [OR] are not obtained, and thus, the LS-category of closed non-manifold Alexandrov $3$-spaces cannot be determined completely using their fundamental groups.
8.B. Using curvature and cohomogeneity
In view of Remark 8.2, computing $\operatorname{\mathsf{dcat}}$ of closed non-manifold Alexandrov $3$-spaces is even more difficult. In fact, $\operatorname{\mathsf{dcat}}$ has not been determined completely even in the manifold case (see Section 5.A). So, in this section, using some well-known results in dimension $3$, we describe up to homeomorphism closed non-manifold Alexandrov $3$-spaces $X$ for which $\operatorname{\mathsf{cat}}(X)>1$ and $\operatorname{\mathsf{dcat}}(X)>1$ when $X$ is non-negatively curved or when a non-trivial compact Lie group acts nicely on $X$.

### v3

@@ 7. For Alexandrov spaces
We now study A. D. Alexandrov spaces with curvature bounded below. In this section, we will extend some of our results for closed manifolds from the previous sections to a more general class of closed Alexandrov spaces, thereby initiating a study of the LS-category of Alexandrov spaces.
7.1. Some definitions
An Alexandrov space is a finite-dimensional, complete, locally compact length space with an inner length metric that locally satisfies a lower curvature bound in the sense of distance comparison. Alexandrov spaces are locally contractible metric ANR, [Wu]. We refer to [BGP][BBI][GGNZ, Section 2], and [Mit, Section 6] for their precise definitions. Examples of Alexandrov spaces include orbit spaces of isometric compact Lie group actions on complete Riemannian manifolds whose sectional curvatures are uniformly bounded below. A non-manifold and non-orbifold example is the suspension of ${\mathbb{C}}P^{2}$, denoted $\operatorname{Susp}({\mathbb{C}}P^{2})$. For more examples and constructions, we refer the reader to [GGNZ, Section 2.3].
@@ 7.1. Some definitions
The connected sum of $M$ and $N$ along $(p,q)$, denoted $M\#^{(p,q)}N$, is defined as the quotient space
$((M\setminus U)\sqcup(N\setminus V))/\sim,$
where $x\sim\psi(x)$ for all $x\in\partial(M\setminus U)$.
Indeed, $M\#^{(p,q)}N$ is a closed Alexandrov $n$-space with curvature bounded below — see, for example, [FGGGH, Section 3] and [GGNZ, Page 9]). Unlike the case of the connected sum of closed manifolds, in general, the sum $M\#^{(p_{1},q_{1})}N$ need not be homeomorphic to $M\#^{(p_{2},q_{2})}N$, even if $M=N$, $p_{1}=q_{1}$, $p_{2}=q_{2}$, and $\varSigma_{p_{1}}=\varSigma_{p_{2}}$, as shown in [FGGGH, Theorem A] for $n=3$. However, when $\varSigma_{p}=\varSigma_{q}=S^{n-1}$ and there is no risk of confusion regarding $p$ and $q$, we will write $M\#N$ to denote the usual connected sum of the Alexandrov spaces $M$ and $N$ along closed $n$-discs.
7.2. Extension of some results
@@
Unlike the case of closed manifolds, the existence of fundamental class and the satisfaction of Poincaré duality is very restrictive in the case of closed Alexandrov spaces, so we will be careful while extending our results.
Let $M$ and $N$ be two closed Alexandrov $n$-spaces such that $N$ is a manifold. We can choose regular points $p\in M$ and $q\in N$ (both $M$ and $N$ have infinitely many regular points) to form $M\#^{(p,q)}N$, which we will write as $M\#N$ from here onwards. Let ${\mathcal{O}}_{N}$ be the orientation sheaf on $N$. We follow the conventions from Section 4. The fundamental class $[M\#N]_{R}$ always exists for either $R={\mathbb{Z}}$ or $R={\mathbb{Z}}_{2}$ (a ${\mathbb{Z}}$-fundamental class exists when both $M$ and $N$ are ${\mathbb{Z}}$-orientable) and gets mapped to $[N]$ under the homomorphism induced in homology by the map $\phi:M\#N\to N$ that collapses $M$ to a point. So, Theorem F, which produces some new computations for the classical LS-category ($\operatorname{\mathsf{cat}}$), can now be proved.
Proof sketch of Theorem F.
The proof for $\operatorname{\mathsf{dcat}}(M\#N)\geq n$ is the same as that of Corollary 4.2 and uses Poincaré duality in the manifold $N$ in ${\mathcal{O}}_{N}\otimes{\mathcal{F}}$ coefficients, where ${\mathcal{F}}$ is a $\pi_{1}(N)$-module such that $\kappa_{n}\in H^{n}(N;{\mathcal{F}})$ is the non-zero primary obstruction to a section of ${\mathcal{B}}_{n}(p_{N}):P(N)_{n}\to N$ due to the asphericity of $N$. Since any closed Alexandrov $n$-space $X$ is an ANR, in particular, locally contractible [Wu], we have $\operatorname{\mathsf{cat}}(X)\leq n$ by [CLOT, Theorem 1.7]. Therefore, $\operatorname{\mathsf{dcat}}(M\#N)=\operatorname{\mathsf{cat}}(M\#N)=n$ is obtained. ∎
@@
Any flat $n$-manifold $N$ is an aspherical Alexandrov $n$-manifold. So, $\operatorname{\mathsf{dcat}}(M\#N)=n$ by Theorem F for any Alexandrov $n$-space $M$. This gives more examples. For $n\geq 4$, let $P^{n-3}$ be an Alexandrov $(n-3)$-space. For any $t\geq 1$, the product #$\#_{i=1}^{t}\operatorname{Susp}({\mathbb{R}}P^{2})\times P^{n-3}$ is a non-manifold Alexandrov $n$-space. So, for any flat $n$-manifold $N$, we get
$\operatorname{\mathsf{dcat}}\left(\left(\#_{i=1}^{t}\operatorname{Susp}({\mathbb{R}}P^{2})\times P^{n-3}\right)\#N\right)=n$
by Theorem F. In particular, we can take $N=\mathcal{N}^{k}\times T^{m}$ for $n=k+m$, where $T^{m}$ is the $m$-torus and $\mathcal{N}^{k}$ is the generalized Hantzsche–Wendt $k$-manifold discussed in Example 4.8. Furthermore, Theorem F also recovers the new computation $\operatorname{\mathsf{dcat}}((\mathcal{N}^{k}\times T^{m})\#(\mathcal{N}^{k}\times T^{m}))=k+m$ obtained in Example 4.8.
In the orientable (or ${\mathbb{Z}}$-orientable) case, a proof simpler than that of Theorem B and Corollary 4.2 can be given for (possibly non-manifold) Alexandrov spaces.
Proof of Theorem G.
Since $N$ is a closed orientable aspherical Alexandrov $n$-space, we can use Theorems 2.3 and 7.2 and [Mit, Theorem 1.8] to get the equality $\operatorname{\mathsf{dcat}}(N)=\operatorname{{\rm cd}}(\pi_{1}(N))=n$. So, in view of Proposition 2.4, the primary obstruction $\kappa_{n}\in H^{n}(N;{\mathcal{F}})$ to a section of the fibration ${\mathcal{B}}_{n}(p_{N}):P(N)_{n}\to N$ is non-zero. Due to orientability, $[M\#N]$ and $[N]$ exist such that $\phi_{*}([M\#N])=[N]$ for the collapsing map $\phi:M\#N\to N$. Clearly, $\phi$ induces an isomorphism of the zeroth homology groups in each coefficient system. Therefore, we have
@@
where $\phi^{*}(\kappa_{n})\in H^{n}(M\#N;\phi^{*}({\mathcal{F}}))$ is the primary obstruction to a lift of $\phi$ with respect to ${\mathcal{B}}_{n}(p_{N})$. Since $\dim(M\#N)=n$, it is the only obstruction to such a lift of $\phi$. Here, ${\mathcal{F}}$ is a $\pi_{1}(N)$-module, hence an abelian group. By Poincaré duality in $N$ in ${\mathcal{F}}$ coefficients (see Theorem 7.2), we have $[N]\frown\kappa_{n}\neq 0$. This means $\phi^{*}(\kappa_{n})\neq 0$. So, $\phi$ does not admit a lift with respect to ${\mathcal{B}}_{n}(p_{N})$. Thus, we get
$n\leq\operatorname{\mathsf{dcat}}(\phi)\leq\operatorname{\mathsf{dcat}}(M\#N)\leq\operatorname{\mathsf{cat}}(M\#N)\leq n$
in view of Propositions 2.5 (2) and 2.6. ∎
@@
We note that the orientability assumption is important to prove Theorem G. If $N$ is not orientable in the sense of [HS], then we will be dealing with ${\mathcal{O}}_{N}\otimes{\mathcal{F}}$ coefficients in $N$ for the cap product $[N]\frown\kappa_{n}$. Here, ${\mathcal{O}}_{N}$ is the “orientation bundle” on $N$, see [Mit, Section 5.3]. So, in this case, while we can still proceed along the lines of Theorem B, we cannot apply Theorem 7.2 to get $[N]\frown\kappa_{n}\neq 0$ because in general, ${\mathcal{O}}_{N}\otimes{\mathcal{F}}\neq{\mathcal{F}}$, even if ${\mathcal{O}}_{N}={\mathbb{Z}}_{2}$.
@@
Following the notations of Example 7.3, we note that Theorem G also recovers the new computation $\operatorname{\mathsf{dcat}}((\mathcal{N}^{k}\times T^{m})\#(\mathcal{N}^{k}\times T^{m}))=k+m$ obtained in Example 4.8 since $\mathcal{N}^{k}$ is orientable for each odd $k\geq 3$.
Theorems F and G both partially extend Corollary 4.2 to some non-manifold cases since they include the possibility of $M$ not being a manifold. In Theorem F, we do not require the orientability of $N$ or $M\#N$ but need $N$ to be a manifold. In Theorem G, we need orientability but do not require any space to be a manifold.
Formally, Theorem G gives new computations of $\operatorname{\mathsf{cat}}$ and $\operatorname{\mathsf{dcat}}$ in orientable non-manifold cases. Unfortunately, we do not have an example of a closed aspherical Alexandrov $n$-space that is not a manifold. For $n=3$, it has been conjectured in [GGNZ, Conjecture 5.1] that such examples do not exist.
@@ 8. For non-manifold Alexandrov $3$-spaces
We now focus on three-dimensional Alexandrov spaces that have recently been studied extensively in [GGG][GGS][NZ], and [FGGGH], for example. In this section, using some well-known results from Alexandrov geometry (in particular, from the above papers), we shall make some observations regarding the classical and the distributional categories of these spaces.
8.1. Initial observations
@@
Given a closed Alexandrov $3$-space $X$, let $M_{X}$ be the compact non-orientable $3$-manifold with boundary obtained by removing disjoint regular neighborhoods of the finitely many singular points of $X$. So, the boundary of $M_{X}$ is a finite collection of ${\mathbb{R}}P^{2}$. Let
$M_{X}=M_{1}\#\cdots\#M_{n}\#l(S^{2}\propto S^{1})$
be the unique prime decomposition [He] of $M_{X}$, where $l\geq 0$ and $S^{2}\propto S^{1}$ denotes the non-orientable $S^{2}$-bundle over $S^{1}$. For each $1\leq i\leq n$, let $\widehat{M_{i}}$ be the Alexandrov $3$-space obtained from $M_{i}$ by capping off its ${\mathbb{R}}P^{2}$ boundary components. Then,
@@
For some $j\leq n$, if $\widehat{M_{j}}$ is orientable and aspherical, or if it is an aspherical manifold, then $\operatorname{\mathsf{dcat}}(X)=3$ in view of Theorems F and G, respectively. The latter situation is possible when $M_{j}$ is aspherical with no ${\mathbb{R}}P^{2}$ boundary components.
@@
We use the above notations. By our assumption, $l\geq 1$. Thus, we can write $X=Y\#(S^{2}\propto S^{1})$, where
$Y=\widehat{M_{1}}\#\cdots\#\widehat{M_{n}}\#(l-1)(S^{2}\propto S^{1}).$
Since $S^{2}\times S^{1}$ is the connected orientable double cover of $S^{2}\propto S^{1}$, the space $Z=Y\#Y\#(S^{2}\times S^{1})$ is a cover of $X$. Since $X$ is a CW complex, so is its cover $Z$. Thus, $(Z,S^{2})$ is a good pair in the sense of [Ha]. So, we can proceed using the Mayer–Vietoris sequence for cohomology of the pair $(Z,S^{2})$ along the lines of [J2, Remark 8.2] to get
$2=c\ell_{{\mathbb{Q}}}(S^{2}\times S^{1})\leq c\ell_{{\mathbb{Q}}}(Z)\leq\operatorname{\mathsf{dcat}}(Z).$
Finally, we use Theorem 2.7 to conclude that $2\leq\operatorname{\mathsf{dcat}}(Z)\leq\operatorname{\mathsf{dcat}}(X)$. ∎
@@
Due to these reasons, the subsequent results in [GLGA] and [OR] are not obtained, and thus, the LS-category of closed non-manifold Alexandrov $3$-spaces cannot be determined completely using just their fundamental groups.
8.2. Using curvature and cohomogeneity
In view of Remark 8.2, computing $\operatorname{\mathsf{dcat}}$ of closed non-manifold Alexandrov $3$-spaces is even more difficult. In fact, $\operatorname{\mathsf{dcat}}$ has not been determined completely even in the manifold case (see Section 5.1). So, in this section, using some well-known results in dimension $3$, we describe up to homeomorphism closed non-manifold Alexandrov $3$-spaces $X$ for which $\operatorname{\mathsf{cat}}(X)>1$ and $\operatorname{\mathsf{dcat}}(X)>1$ when $X$ is non-negatively curved or when a non-trivial compact Lie group acts nicely on $X$.